# Non-abelian Hodge correspondence over singular Kähler spaces

### v1

@@ 3.3.2. Approximate Hermitian-Einstein structure
$=$ $\int_{Y_{\mathrm{orb}}\setminus f_{\mathrm{orb}}^{-1}(\Sigma)}\operatorname{tr}(\pi_{\alpha_{\mathrm{orb}}}\sqrt{-1}\Lambda_{f_{\mathrm{orb}}^{*}\omega}F_{\hat{H}_{\mathrm{orb}},\theta_{\mathrm{orb}}})-|\overline{\partial}_{\theta_{\mathrm{orb}}}\pi_{\alpha,\mathrm{orb}}|_{\hat{H}_{\mathrm{orb}},f^{*}\omega_{\mathrm{orb}}}^{2}\frac{f^{*}\omega_{\mathrm{orb}}^{n}}{n!}$
@@
$=$ $\int_{X\setminus\Sigma}\operatorname{tr}(\pi_{\alpha}\sqrt{-1}\Lambda_{\omega}F_{\hat{H},\theta_{X_{\mathrm{reg}}}})-|\overline{\partial}_{\theta_{X_{\mathrm{reg}}}}\pi_{\alpha}|_{\hat{H},\omega}^{2}\frac{\omega^{n}}{n!}$
@@
where $\pi_{\alpha,\mathrm{orb}}$ denotes the weakly holomorphic subbundle associated to $\mathcal{E}_{\alpha,\mathrm{orb}}$. Furthermore, by the argument in Simpson’s paper (page 888 in [97]) and the inequality (3.62), we have
@@
(3.63) $\begin{split}\int_{X\setminus\Sigma}\operatorname{tr}(u_{\infty}\sqrt{-1}\Lambda_{\omega}F_{\hat{H},\theta})+\langle\sum_{\alpha=1}^{l-1}(\mu_{\alpha+1}-\mu_{\alpha})(dP_{\alpha})^{2}(u_{\infty})(\overline{\partial}_{\theta}u_{\infty}),\overline{\partial}_{\theta}u_{\infty}\rangle_{\hat{H}}\leq-r^{-\frac{1}{2}}\frac{C^{\ast}}{\hat{C}_{1}}.\end{split}$
This complete the proof of (ii). ∎
@@
Let $H(t)$ be the long time solution of (3.9) on the Higgs bundle $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{\mathrm{reg}})|_{X\setminus\Sigma}$ with initial metric $\hat{H}$. By Proposition 3.22 (ii), if the Higgs sheaf $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ is $(\omega_{0},\cdots,\omega_{n-2})$-stable, then $\|\log(H^{-1}(1)H(t))\|_{L^{1}}$ are uniformly bounded, i.e. we obtain uniform $C^{0}$-estimates on $X\setminus\Sigma$. By the standard parabolic estimates, we can get uniform local $C^{\infty}$-estimates, and then, by choosing a subsequence, $H(t)\rightarrow H_{\infty}$ in $C^{\infty}_{loc}(X\setminus\Sigma)$-topology, and (3.42) implies
(3.64) $\sqrt{-1}\Lambda_{\omega}(F_{H_{\infty}}+[\theta_{X_{\mathrm{reg}}},(\theta_{X_{\mathrm{reg}}})^{*H_{\infty}}])=\lambda\cdot\mathrm{Id}_{E}.$
Using Proposition 3.22 again, we known that, if the Higgs bundle $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})|_{X\setminus\Sigma}$ is $(\omega_{0},\cdots,\omega_{n-2})$-semistable, then
@@
Combining (3.15), we obtain (3.11). This completes the proof of Theorem 3.5.
4. Characterization of harmonic bundles over the regular loci
@@ 4.1. Basic facts
  1. (4.1.1)
    For any locally free Higgs sheaf $(\mathcal{E}_{X},\theta_{X})$ such that $\pi^{*}(\mathcal{E}_{X},\theta_{X})$ is semistable with (1.1) for some single Kähler form $\omega_{\widehat{X}}$, then $(\mathcal{E}_{X},\theta_{X})$ is semi-stable with respect to any nef forms $\eta_{0},\cdots,\eta_{n-2}$ and all Chern classes $c_{i}(\mathcal{E}_{X})\in H^{2i}(X,\mathbb{Q})$ vanish.
  2. (4.1.2)
@@
The item (4.1.1) follows from the same proof of [38, Lemma 7.8]. The item (4.1.2) follows from Takayama’s result [101] that the induced morphism of $\pi_{*}:\pi_{1}(\widehat{X})\rightarrow\pi_{1}(X)$ is an isomorphism, the proof is purely analytic and thus also holds for analytic klt spaces. ∎
Let us follow [34, 15, 55, 28] to review the concepts of maximally quasi-étale cover, which takes a central role in our construction.
@@
A normal complex space $X$ is said to be maximally quasi-étale if $X$ satisfies the following equivalent conditions:
@@
    Any étale cover of $X_{\mathrm{reg}}$ can be extended to a étale cover of $X$.
@@
Remark 4.3 (The extension of linear representations).
If $X$ is maximally quasi-étale, then it follows from Malcev’s theorem that any linear representation of $\pi_{1}(X_{\mathrm{reg}})$ factors through $\pi_{1}(X)$ (cf. [36, Proof of Proposition 3.10]), i.e.
@@
Since $i_{*}:\pi_{1}(X_{\mathrm{reg}})\rightarrow\pi_{1}(X)$ is surjective (cf. [65, Proposition 2.10]), $\rho_{0}$ is semi-simple if and only if $\rho$ is semi-simple. This is equivalent to that, any (resp. semi-simple) local system on $X\setminus\Sigma$ for some analytic subspace of $\mathrm{codim}_{Z}X\geq 2$, can be extended to a (resp. semi-simple) local system on $X$ by combining $\pi_{1}(X_{\mathrm{reg}}\setminus\Sigma)\cong\pi_{1}(X_{\mathrm{reg}})$.
Proposition 4.4 (The existence of maximally quasi-étale covers).
@@
The existence of a maximally quasi-étale cover of a projective klt variety was proved by [34]. Building on the recent work of Fujino [29] of relative MMP in the context of analytic spaces, the existence of maximally quasi-étale covers for analytic klt spaces can be obtained by following the same method in [34] (see e.g. [28, Theorem 2.4 and Theorem 2.5]).
@@
Let $X$ be a compact Kähler klt space and $\gamma:Y\rightarrow X$ be a maximally quasi-étale cover of $X$ induced by some $G\subset\mathrm{Aut}(Y)$. Then the following hold:
  1. (4.6.1)
@@
  4. (4.6.4)
    For any $E_{X_{\mathrm{reg}}}\in\mathrm{LSys}_{X_{\mathrm{reg}}}$ (resp. $\mathrm{sLSys}$), $(\gamma|_{\gamma^{-1}(X_{\mathrm{reg}})})^{*}E_{X_{\mathrm{reg}}}$ extends to a $G$-equivariant local system $\rho_{Y}\in\mathrm{LSys}_{Y}$ (resp. $\mathrm{sLSys}_{Y}$). Conversely, the restriction of a $G$-equivariant local system $\rho_{Y}\in\mathrm{LSys}_{Y}$ (resp. $\mathrm{sLSys}_{Y}$) descends to a local system $\rho_{X_{\mathrm{reg}}}\in\mathrm{sLSys}_{X_{\mathrm{reg}}}$.
@@
The item (4.6.1) can be refered to [31, Proposition 3.6]. The item (4.6.2) follows from the purity of branch theorem due to Grauert-Remmert [32]. The item (4.6.3) follows from Lemma 2.21, Lemma 2.24 and the item (4.6.1). The item (4.6.4) follows from Remark 4.3. It also hold for semi-simple local systems bacause $\gamma_{*}:\pi_{1}(\gamma^{-1}(X_{\mathrm{reg}}))\subset\pi_{1}(Y_{\mathrm{reg}})$ is injective by the item (4.6.2). ∎
@@ 4.2. Harmonic bundles and descent result
This section is a study on harmonic bundle structure over the regular locus of a compact Kähler klt space $X$. The following standard definition is given to fix the conventions (cf. [106, Section 2]).
Let $(E,D)$ be a flat bundle equipped with a Hermitian metric $H$ on a complex complex manifold $M$. There exists a unique decomposition of $D$ by
(4.1) $D=D_{H}+\Psi_{H}$
where $D_{H}$ is a $H$-connection and $\Psi_{H}\in\Lambda^{1}(M,\operatorname{End}(E))$ is self-adjoint with respect to $H$. We define
@@
We say that $(E,D,H)$ is a harmonic bundle or a flat bundle $(E,D)$ admits a pluri-harmonic metric $H$ if $G_{H}=0$, which is equivalent to that
(4.3) $(D_{H}^{0,1})^{2}=0,\ D_{H}^{0,1}\Psi_{H}^{1,0}=0\text{ and }\Psi_{H}^{1,0}\wedge\psi_{H}^{1,0}=0.$
Thus, $(E,D,H)$ is a harmonic bundle if and only if $(E,\bar{\partial},\theta):=(E,D_{H}^{0,1},\Psi_{H}^{1,0})$ is a Higgs bundle and we could write the decomposition (4.1) as
@@ 4.2.1. Pluriharmonic metrics on Harmonic bundles
This subsubsection is devoted to prove the following statement.
@@
Let $X$ be a compact Kähler klt space and $Z$ be an analytic subspace of $X$ with $\mathrm{codim}_{Z}X\geq 2$. Suppose that $X\setminus Z$ is smooth. Then the following hold:
@@
    A flat bundle on $X\setminus Z$ is semi-simple if and only if it admits a pluri-harmonic metric, which is unique up to some constant on each simple component.
@@
Because $i_{*}:\pi_{1}(X_{\mathrm{reg}}\setminus Z)\rightarrow\pi_{1}(X_{\mathrm{reg}})$ is an isomorphism, any semi-simple flat bundle on $X\setminus Z$ could extends to a semi-simple flat bundle on $X_{\mathrm{reg}}$ and then (2) is a direct consequence of (1).
@@
It remains to verify the uniqueness and that the existence of a pluri-harmonic metric implies semi-simpleness. The classical proof in the non-compact context via the integral method is not available because we do not put any assumption such as “admissibility” on pluri-harmonic metrics. We adapt another method via the maximum principle, which is well‑adapted to our setting thanks to Grauert–Remmert’s extension theorem [32].
@@
Let $(E,D,H)$ be a harmonic bundle on a complex manifold $M=X\setminus Z$ for some compact analytic space $X$ and an analytic subspace $Z$ of codimension at least $2$. Then for any $D$-invariant subbundle $W$ of $E$ such that $(W,D_{W})$ is simple where $D_{W}=D|_{W}$, it holds that
@@
$D$ takes the form
(4.6) $D=\left(\begin{matrix}D_{W}&\beta_{W}\\ 0&D_{W^{\perp}}\end{matrix}\right)$
where $\beta_{W}\in\Lambda^{1}(M,\operatorname{Hom}(W^{\perp},W))$. Recall (4.4) that $D$ uniquely decomposes as $D_{H}+\Psi_{H}$. It is easily known that $D_{H}$ and $\Psi_{H}$ take the form
(4.7) $D_{H}=\left(\begin{matrix}D_{H_{W}}&\frac{1}{2}\beta_{W}\\ -\frac{1}{2}\beta^{*}_{W}&D_{H_{W^{\perp}}}\end{matrix}\right)\ \text{ and }\ \Psi_{H}=\left(\begin{matrix}\Psi_{H_{W}}&\frac{1}{2}\beta_{W}\\ \frac{1}{2}\beta^{*}_{W}&\Psi_{H_{W^{\perp}}}\end{matrix}\right)$
where we decompose $D_{W}=D_{H_{W}}+\Psi_{H_{W}}$ and $D_{W^{\perp}}=D_{H_{W^{\perp}}}+\Psi_{H_{W^{\perp}}}$ (cf. [106, Section 2.3]). Recall the pluri-harmonicity of $H$ means (4.3). Combining this with (4.7) gives
@@
Recall the simple flat bundle $(W,D_{W})$ admits a pluri-harmonic metric $K$ as explained above. Then $D_{K}^{0,1}\Psi_{K}^{1,0}=0$ and (4.8) imply
@@
where $h=K^{-1}H_{W}$ (cf. [93, (2.25)] for the derivation of the first equality). Consequently, $\log\det(h)$ is a plurisubharmonic function on $M$ and so can be extended to a plurisubharmonic function on $X$ due to the extension theorem by Grauert-Remmert [32, page 181] since $\mathrm{codim}_{Z}X\geq 2$. Thus, $\log\det(h)$ must be constant. Again using (4.9), we deduce that $\beta_{W}^{1,0}=\beta_{W}^{0,1}=0$ and so $\beta_{W}=0$, which completes the proof. ∎
To prove the uniqueness, it suffices to consider the case when $(E,D)$ is simple by Lemma 4.10. Suppose that $H,K$ are two pluri-harmonic metrics on $(E,D)$, then we have
(4.10) $\sqrt{-1}\partial\bar{\partial}\operatorname{tr}(h)=\frac{1}{2}\sqrt{-1}\{Dh\cdot h^{-\frac{1}{2}},\overline{Dh\cdot h^{-\frac{1}{2}}}\}_{K}+\operatorname{tr}(h\sqrt{-1}(G_{H}-G_{K})),$
where $h=K^{-1}H\in C^{\infty}(M,\operatorname{End}(E))$ (cf. [93, (2.36)]). Consequently, $\operatorname{tr}(h)$ is a global plurisubharmonic function on $M$ because $G_{H}=0$ and $G_{K}=0$ (cf. (4.3)) and so must be some constant as above. Applying (4.10) again gives $Dh=0$ over $M$. Then $h=c\operatorname{Id}_{E}$ for some constant $c>0$. Otherwise, $(E,D)$ must split according to the eigenspaces of $h$. Therefore $H=cK$.
It remains to show that the existence of a pluri-harmonic metric $H$ implies the semi-simpleness of $(V,D)$, which can be concluded by Lemma 4.10 and an induction argument because any flat bundle $(E,D)$ includes a $D$-invariant simple subbundle. Thus, the proof of Proposition 4.8 is complete. ∎

### v2

@@ 3.3.2. Approximate Hermitian-Einstein structure
$=$ $\int_{Y_{\mathrm{orb}}\setminus f_{\mathrm{orb}}^{-1}(\Sigma)}\operatorname{tr}(\pi_{k,\mathrm{orb}}\sqrt{-1}\Lambda_{f_{\mathrm{orb}}^{*}\omega}F_{\hat{H}_{\mathrm{orb}},\theta_{\mathrm{orb}}})-|\overline{\partial}_{\theta_{\mathrm{orb}}}\pi_{k,\mathrm{orb}}|_{\hat{H}_{\mathrm{orb}},f_{\mathrm{orb}}^{*}\omega}^{2}\frac{(f_{\mathrm{orb}}^{*}\omega)^{n}}{n!}$
@@
$=$ $\int_{X\setminus\Sigma}\bigg(\operatorname{tr}(\pi_{k}\sqrt{-1}\Lambda_{\omega}F_{\hat{H},\theta_{X_{\mathrm{reg}}}})-|\overline{\partial}_{\theta_{X_{\mathrm{reg}}}}\pi_{k}|_{\hat{H},\omega}^{2}\bigg)\frac{\omega^{n}}{n!},$
@@
where $\pi_{k,\mathrm{orb}}$ denotes the weakly holomorphic subbundle associated with $\mathcal{E}_{k,\mathrm{orb}}$. Furthermore, applying the argument in Simpson’s paper [100, Page 888] and the inequality (3.62), we conclude
@@
(3.63) $\begin{split}&\int_{X\setminus\Sigma}\bigg(\operatorname{tr}(u_{\infty}\sqrt{-1}\Lambda_{\omega}F_{\hat{H},\theta})+\langle\sum_{k=1}^{l-1}(\mu_{k+1}-\mu_{k})(dP_{k})^{2}(u_{\infty})(\overline{\partial}_{\theta}u_{\infty}),\overline{\partial}_{\theta}u_{\infty}\rangle_{\hat{H}}\bigg)\frac{\omega^{n}}{n!}\\ &\leq-r^{-\frac{1}{2}}\frac{C^{\ast}}{\hat{C}_{1}^{\prime}}.\end{split}$
This completes the proof of (3.49). ∎
@@
Let $H(t)$ be the long time solution of (3.9) on the Higgs bundle $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{\mathrm{reg}})|_{X\setminus\Sigma}$ with initial metric $\hat{H}$. By Proposition 3.23, if the Higgs sheaf $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ is $(\omega_{0},\cdots,\omega_{n-2})$-stable, then $\|\log(H^{-1}(1)H(t))\|_{L^{1}(X\setminus\Sigma,H(1))}$ are uniformly bounded, i.e., we obtain uniform $C^{0}$-estimates on $X\setminus\Sigma$. From the standard parabolic estimates, we can get uniform local $C^{\infty}$-estimates. Then after passing to a subsequence, $H(t)\rightarrow H_{\infty}$ in $C^{\infty}_{loc}(X\setminus\Sigma)$-topology, and (3.42) implies
(3.64) $\sqrt{-1}\Lambda_{\omega}(F_{H_{\infty}}+[\theta_{X_{\mathrm{reg}}},(\theta_{X_{\mathrm{reg}}})^{*H_{\infty}}])=\lambda\cdot\mathrm{Id}_{\mathcal{E}_{X_{\mathrm{reg}}}}$
on $X\setminus\Sigma$. Using Proposition 3.23 again, we know that if the Higgs bundle $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})|_{X\setminus\Sigma}$ is $(\omega_{0},\cdots,\omega_{n-2})$-semistable, then
@@
Combining this with (3.15), we obtain (3.11). This completes the proof of Theorem 3.5.
4. Characterization of harmonic bundles over the regular locus
@@ 4.1. Basic facts
  1. (4.1.a)
    For any locally free Higgs sheaf $(\mathcal{E}_{X},\theta_{X})$ such that $\pi^{*}(\mathcal{E}_{X},\theta_{X})$ is semistable and satisfies (1.1) for some Kähler form $\omega_{\widehat{X}}$, then $(\mathcal{E}_{X},\theta_{X})$ is semistable with respect to any nef forms $\eta_{0},\cdots,\eta_{n-2}$ and all Chern classes $c_{i}(\mathcal{E}_{X})\in H^{2i}(X,\mathbb{Q})$ vanish.
  2. (4.1.b)
@@
Item (4.1.a) follows from the same proof as [38, Lemma 7.8]. Item (4.1.b) follows from Takayama’s result [104], which states that the induced morphism $\pi_{*}:\pi_{1}(\widehat{X})\rightarrow\pi_{1}(X)$ is an isomorphism. The proof is purely analytic and thus also holds for analytic klt spaces. ∎
Let us follow [34, 15, 56, 28] to review the concept of a maximally quasi-étale cover, which plays a central role in our construction.
@@
A normal complex space $X$ is said to be maximally quasi-étale if it satisfies the following equivalent conditions:
@@
    Any étale cover of $X_{\mathrm{reg}}$ extends to an étale cover of $X$.
@@
Remark 4.3 (Extension of linear representations).
If $X$ is maximally quasi-étale, then Malcev’s theorem implies that any linear representation of $\pi_{1}(X_{\mathrm{reg}})$ factors through $\pi_{1}(X)$ (see e.g. the proof of [36, Proposition 3.10]), i.e.,
@@
Since $i_{*}:\pi_{1}(X_{\mathrm{reg}})\rightarrow\pi_{1}(X)$ is surjective (cf. [66, Proposition 2.10]), $\rho_{0}$ is semisimple if and only if $\rho$ is semisimple. This implies that any (resp. semisimple) local system on $X\setminus\Sigma$, for some analytic subspace $\Sigma\subset X$ with $\mathrm{codim}_{X}\Sigma\geq 2$, extends to a (resp. semisimple) local system on $X$, based on the standard fact that $\pi_{1}(X_{\mathrm{reg}}\setminus\Sigma)\cong\pi_{1}(X_{\mathrm{reg}})$.
Theorem 4.4 (The existence of maximally quasi-étale covers, see e.g. [28, Theorems 2.4 and 2.5]).
@@
The existence of a maximally quasi-étale cover for a projective klt variety was proved in [34]. Building on Fujino’s recent work [29] on the relative MMP for projective morphisms between analytic spaces, the existence of maximally quasi-étale covers for analytic klt spaces can be obtained by following the same method as in [34].
@@
Let $X$ be a compact Kähler klt space and $\gamma:Y\rightarrow X$ be a maximally quasi-étale cover induced by some subgroup $G\subset\mathrm{Aut}(Y)$. Then the following hold:
  1. (4.6.a)
@@
  4. (4.6.d)
    For any $E_{X_{\mathrm{reg}}}\in\mathrm{LSys}_{X_{\mathrm{reg}}}$ (resp. $\mathrm{sLSys}_{X_{\mathrm{reg}}}$), the pullback $(\gamma|_{\gamma^{-1}(X_{\mathrm{reg}})})^{*}E_{X_{\mathrm{reg}}}$ extends to a $G$-equivariant local system $\rho_{Y}\in\mathrm{LSys}_{Y}$ (resp. $\mathrm{sLSys}_{Y}$). Conversely, the restriction of a $G$-equivariant local system $\rho_{Y}\in\mathrm{LSys}_{Y}$ (resp. $\mathrm{sLSys}_{Y}$) descends to a local system $\rho_{X_{\mathrm{reg}}}\in\mathrm{sLSys}_{X_{\mathrm{reg}}}$.
@@
Item (4.6.a) can be refered to [31, Proposition 3.6]. Item (4.6.b) is a consequence of the purity of branch theorem due to Grauert-Remmert [32]. Item (4.6.c) follows from Proposition 2.21, Lemma 2.24 and item (4.6.a). Item (4.6.d) follows from Remark 4.3. The statement also holds for semisimple local systems because $\gamma_{*}:\pi_{1}(\gamma^{-1}(X_{\mathrm{reg}}))\subset\pi_{1}(Y_{\mathrm{reg}})$ is injective by item (4.6.b). ∎
@@ 4.2. Harmonic bundles and descent result
This subsection studies harmonic bundle structures over the regular locus of a compact Kähler klt space. The following standard definition fixs the conventions (see [109, Section 2] for general Riemannian setting).
Let $(E,D)$ be a flat bundle equipped with a Hermitian metric $H$ on a complex manifold $M$. There exists a unique decomposition of $D$ as
(4.1) $D=D_{H}+\Psi_{H},$
where $D_{H}$ is an $H$-connection and $\Psi_{H}\in\Lambda^{1}(M,\operatorname{End}(E))$ is self-adjoint with respect to $H$. We define
@@
We say that $(E,D,H)$ is a harmonic bundle, or that the flat bundle $(E,D)$ admits a pluri-harmonic metric $H$, if $G_{H}=0$. This condition is equivalent to
(4.3) $(D_{H}^{0,1})^{2}=0,\quad D_{H}^{0,1}\Psi_{H}^{1,0}=0\ \text{ and}\ \ \Psi_{H}^{1,0}\wedge\Psi_{H}^{1,0}=0.$
Thus, $(E,D,H)$ is a harmonic bundle if and only if $(E,\bar{\partial},\theta):=(E,D_{H}^{0,1},\Psi_{H}^{1,0})$ is a Higgs bundle and the decomposition (4.1) can be written as
@@ 4.2.1. Pluriharmonic metrics on Harmonic bundles
This subsubsection is devoted to proving the following statement.
@@
Let $X$ be a compact Kähler klt space and $Z\subset X$ be an analytic subspace with $\mathrm{codim}_{X}Z\geq 2$. Suppose that $X\setminus Z$ is smooth. Then the following hold:
@@
    A flat bundle on $X\setminus Z$ is semisimple if and only if it admits a pluri-harmonic metric. Such a metric is unique up to a positive constant on each direct summand.
@@
Because the inclusion induces an isomorphism $i_{*}:\pi_{1}(X_{\mathrm{reg}}\setminus Z)\rightarrow\pi_{1}(X_{\mathrm{reg}})$, any semisimple flat bundle on $X\setminus Z$ extends to a semisimple flat bundle on $X_{\mathrm{reg}}$. Thus (2) is a direct consequence of (1).
@@
It remains to verify the uniqueness and to show that the existence of a pluri-harmonic metric implies semisimplicity. The standard proof in the noncompact Kähler context via the integral method is not available because we do not impose any assumption such as “admissibility” on the pluri-harmonic metrics. We adapt a new argument via the maximum principle, which is well‑suited to our setting thanks to Grauert-Remmert’s extension theorem [32].
@@
Let $(E,D,H)$ be a harmonic bundle on a complex manifold $M=X\setminus Z$ for some compact analytic space $X$ and an analytic subspace $Z$ of codimension at least $2$. Then for any $D$-invariant subbundle $W$ of $E$ such that $(W,D_{W})$ is simple, where $D_{W}=D|_{W}$, it holds that
@@
With respect to the orthogonal decomposition $E=W\oplus W^{\perp}$, the connection $D$ takes the form
(4.6) $D=\left(\begin{matrix}D_{W}&\beta_{W}\\ 0&D_{W^{\perp}}\end{matrix}\right),$
where $\beta_{W}\in\Lambda^{1}(M,\operatorname{Hom}(W^{\perp},W))$. Recall from (4.4) that $D$ admits a unique decomposition $D_{H}+\Psi_{H}$. One can verify that $D_{H}$ and $\Psi_{H}$ are then given by
(4.7) $D_{H}=\left(\begin{matrix}D_{H_{W}}&\frac{1}{2}\beta_{W}\\ -\frac{1}{2}\beta^{*}_{W}&D_{H_{W^{\perp}}}\end{matrix}\right)\ \text{ and }\ \Psi_{H}=\left(\begin{matrix}\Psi_{H_{W}}&\frac{1}{2}\beta_{W}\\ \frac{1}{2}\beta^{*}_{W}&\Psi_{H_{W^{\perp}}}\end{matrix}\right),$
where we have used the decompositions $D_{W}=D_{H_{W}}+\Psi_{H_{W}}$ and $D_{W^{\perp}}=D_{H_{W^{\perp}}}+\Psi_{H_{W^{\perp}}}$ (cf. [109, Section 2.3]). Recall that the pluri-harmonicity of $H$ is equivalent to (4.3). Combining this with (4.7) gives
@@
Recall that the simple flat bundle $(W,D_{W})$ admits a pluri-harmonic metric $K$, as explained above. Then $D_{K}^{0,1}\Psi_{K}^{1,0}=0$, and together with (4.8) this implies
@@
where $h=K^{-1}H_{W}$ (cf. [95, (2.25)] for the derivation of the first equality). Consequently, $\log\det(h)$ is a plurisubharmonic function on $M$ and thus extends to a plurisubharmonic function on $X$ due to Grauert-Remmert’s extension theorem [32, Page 181], since $\mathrm{codim}_{X}Z\geq 2$. Hence, $\log\det(h)$ must be constant on $X$. Using (4.9) again, we deduce that $\beta_{W}^{1,0}=\beta_{W}^{0,1}=0$, and so $\beta_{W}=0$, which completes the proof. ∎
To prove the uniqueness, it suffices to consider the case when $(E,D)$ is simple by Lemma 4.10. Suppose that $H$ and $K$ are two pluri-harmonic metrics on $(E,D)$. Then we have
(4.10) $\sqrt{-1}\partial\bar{\partial}\operatorname{tr}(h)=\frac{1}{2}\sqrt{-1}\langle Dh\cdot h^{-\frac{1}{2}},\overline{Dh\cdot h^{-\frac{1}{2}}}\rangle_{K}+2\operatorname{tr}(h\sqrt{-1}(G_{H}-G_{K})),$
where $h=K^{-1}H\in C^{\infty}(M,\operatorname{End}(E))$ with $M=X\setminus Z$ (cf. [95, Section 2]). Consequently, $\operatorname{tr}(h)$ is a global plurisubharmonic function on $M$ because $G_{H}=0$ and $G_{K}=0$ (cf. (4.3)), and thus, as argued above, must be constant. Applying (4.10) again gives $Dh=0$ over $M$. Then $h=c\operatorname{Id}_{E}$ for some constant $c>0$. Otherwise, $(E,D)$ would split according to the eigenspaces of $h$. Therefore, $H=cK$.
It remains to show that the existence of a pluri-harmonic metric $H$ implies the semisimplicity of $(E,D)$. This can be concluded by Lemma 4.10 and an induction argument, because any flat bundle $(E,D)$ contains a $D$-invariant simple subbundle. Hence the proof of Proposition 4.8 is complete. ∎